A company had $18 of sales per share for the year that just ended. You expect the company to grow their sales by 6.5 percent for the next five years. After that, you expect the company to grow 3.5 percent in perpetuity. The company has a 14 percent ROE and you expect that to continue forever. The company's net margins are 6 percent and the cost of equity is 11 percent. Use the free cash flow to equity model to value this stock.
Answer #1
Year 0 Year 1 Year 2 Year 3 Year 4 Year 5
Sales per share 18
Earnings per share calculation =18*6% =1.08*1.065 =1.15*1.065 =1.30*1.065 =1.39*1.065 =1.39*1.065
Earnings per share 1.08 1.15 1.22 1.30 1.39 1.48
Terminal value =1.48*(1+3.5%)/(11%-3.5%)
Terminal value calculation 20.42
FCFE 1.15 1.22 1.30 1.39 21.90
Present value calculation =1.15/(1+11%)^1 =1.22/(1+11%)^2 =1.3/(1+11%)^3 =1.39/(1+11%)^4 =21.9/(1+11%)^5
Present value 1.04 0.99 0.95 0.92 13.00
Total present value 16.89
Value per share = $ 16.89